# Evaluating Robustness of LLMs in Question Answering on Multilingual Noisy OCR Data

### v2

@@ Abstract.
copyright: acmlicenseddoi: XXXXXXX.XXXXXXXconference: The 34th ACM International Conference on Information and knowledge Management; November 10–14, 2025; Seoul, Koreaisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Question answeringccs: Information systems Content analysis and feature selection
@@ 1. Introduction
Historical corpora have been already used as an underlying text for numerous natural language processing (NLP) tasks, including Named Entity Recognition (NER), topic modeling, text classification, neural ranking, and understanding semantic changes in language over time (Ehrmann et al., 2020, 2023; Yang et al., 2011; Liebeskind and Liebeskind, 2020; Subramani et al., 2020). While these tasks offer valuable insights, it is expected that automatic question answering (QA) systems become an important interface to the vast historical collections. QA systems enable direct and intuitive access to specific information, offering a powerful means of analyzing, searching and understanding historical texts. By retrieving precise answers to user queries, QA can significantly enhance the accessibility, interpretability, and utility of historical corpora, making them more easily accessible and more useful for scholars, researchers, and the general public alike.
Figure 1. An example of CorrectedOCR and RawOCR text from the MultiOCR dataset for the English language, highlighting different types of errors along with questions corresponding to this text. WER and CER denote Word Error Rate and Character Error Rate, respectively, indicating the level of errors in the text. The green highlights represent insertion errors, where IEO denotes Insertion Edit Operations - the number of insertions needed to transform RawOCR into CorrectedOCR. Similarly, red and blue highlights indicate deletion and substitution errors, with DEO and SEO representing Deletion Edit Operations and Substitution Edit Operations, respectively. The black boxes with numbers in the CorrectedOCR and RawOCR text correspond to the answers for each question in the paragraph.
@@
Although extensive research has focused on improving OCR accuracy and post-processing correction techniques, the specific impact of OCR noise on QA remains largely unexplored. Previous studies have examined challenges related to OCR in information retrieval (IR) (de Oliveira et al., 2023), historical text processing (Piotrowski, 2012), and entity recognition (Grover et al., 2008). However, a systematic investigation of how OCR errors affect QA model performance is still missing. Furthermore, our study contributes also to the body of work on robustness of LLMs against noisy inputs such as noisy prompts (Wang et al., 2023), however we do it in a quite novel yet realistic setting.
In this paper, we address this gap in understanding how large language models (LLMs) perform in question-answering (QA) tasks when dealing with noisy OCR-generated text. In particular, we first introduce MultiOCR-QA  [^2] The dataset and code will be made available after paper publication. a new multilingual QA dataset covering English, French, and German. This dataset includes both RawOCR text (OCR-generated text with errors) and CorrectedOCR text (ground truth), allowing for a direct comparison of QA performance under different text quality conditions. To generate contextually relevant question-answer pairs from historical text excerpts, we leverage instruction-fine-tuned LLMs. We then systematically evaluate the impact of different types of OCR errors—insertions, deletions, and substitutions—on QA model performance, offering new insights into the strengths and limitations of LLMs when processing noisy historical data. Figure 1 illustrates an example from the MultiOCR-QA dataset with various OCR errors impacting text accuracy. It presents QA pairs and the corresponding responses for questions when using CorrectedOCR and RawOCR text as input context to Gemma 2–27b, demonstrating the effects of OCR noise on QA performance. For instance, Question 1: ”What did the man say happened to his leg?” produces different answers depending on whether clean or noisy text was input. When using the CorrectedOCR context, the model produces the response: ”He said his sense of the leg was gone.” In contrast, the response based on the RawOCR context is: ”The fence of his leg is gone,” demonstrating how OCR-induced noise can significantly alter the meaning of the answer and degrade the QA accuracy.
@@ 2. Related Work
Optical Character Recognition (OCR) technology has significantly contributed to the digitization of historical documents, enabling large-scale access to printed and handwritten archives. However, the accuracy of OCR output is highly dependent on factors such as document quality, font variations, language complexity, and historical text degradation. Early studies on OCR error analysis focused primarily on measuring character- and word-level error rates to assess their impact on text-based applications (Piotrowski, 2012; Hill and Hengchen, 2019).
@@ 3.1. Data Collection
  • French: For the French language, the ICDAR 2019 dataset provides a collection obtained from three sources: the HIMANIS [^6] https://www.himanis.org Project, IMPACT - National Library of France, and the RECEIPT [^7] http://findit.univ-lr.fr/ dataset. The original dataset contained 2,800 files.
@@ 3.2. Question-Answer Generation
For fine-tuning the model in English, we used the SQuAD v1 dataset (Rajpurkar et al., 2016). We randomly selected 2,067 paragraphs and 10,570 questions from the development set and 3,000 paragraphs and 13,894 questions from the test set. Similarly, for the French language, we fine-tuned the model on the FQuAD dataset (d’Hoffschmidt et al., 2020), utilizing both the validation and training sets, comprising 5,689 paragraphs and 23,919 questions. For the German language, we fine-tuned the model on the GermanQuAD dataset (Möller et al., 2021) using the training and validation sets, which consisted of 3,014 paragraphs and 13,722 questions. This fine-tuning step ensured that the model accurately generated factoid-style QA pairs, reducing instances of open-ended or conversational questions.
After instruction fine-tuning, we used the fine-tuned model for each language to generate questions for the preprocessed dataset prepared in the initial step. To systematically generate high-quality QA pairs, we employed a two-step prompt-based approach: Answer Extraction and Question Generation for the Extracted Answers.
@@
Dataset Filtering: After generating the dataset for each language, we applied additional filtering steps to ensure quality and consistency. Specifically, we removed questions that did not end with a question mark, duplicate questions, and questions with excessively long answers. Since LLM-generated datasets may contain hallucinated long answers, we applied this additional filtering by removing excessively long answers.
As a result, we removed 66 questions for English Language, 530 questions for French Language, and 7,210 questions for German Language. This final filtering step ensured that the MultiOCR-QA dataset consisted of concise, well-structured, and factually accurate question-answer pairs.
@@ 4. Dataset Analysis
After applying all the filtering steps, we obtained the final dataset, comprising 50,079 question-answer pairs.
The dataset statistics, including average paragraph length, question length, and answer length, are presented in Table 1.
@@
Avg. CorrectedOCR paragraph length (words) 271.73 297.53 212.86
@@
Avg. answer length (words) 2.05 3.12 5.63
@@
4.1. Quantifying and Filtering Noise
To assess the impact of OCR noise on QA quality, we quantified the noise level in the RawOCR text using two standard metrics: Character Error Rate (CER) and Word Error Rate (WER). CER measures the proportion of character-level errors in the raw OCR text compared to the ground truth text. It is computed as the number of insertions, deletions, and substitutions (including spaces) required to transform the RawOCR text into its correct form. WER quantifies word-level discrepancies, representing the proportion of words that require modifications (insertions, deletions, or substitutions) to match ground truth text. Both CER and WER were computed using the Levenshtein distance (Miller et al., 2009), which determines the minimum number of edits needed to correct the OCR-generated text. A high CER but low WER suggests that errors are concentrated within a few words (e.g., spelling variations), whereas a high WER indicates distortions across multiple words, significantly affecting readability.
Outlier Detection: To ensure a reliable analysis, we applied the Interquartile Range (IQR) method to detect and remove outliers in CER values. Outliers were defined as CER values below Q1 - 1.5 * IQR or above Q3 + 1.5 * IQR, where Q1 and Q3 represent the 33th and 66th percentiles of the CER distribution, respectively, and IQR is the difference between Q3 and Q1. This filtering resulted in the removal of 25 English, 351 French, and 2,423 German paragraphs.
@@ 5.1. Experimental Settings
We conducted experiments using multiple large language models (LLMs), including Qwen2.5 7B (Yang et al., 2024), LLaMa 3.1-8B (Dubey et al., 2024), Gemma-2-27B (Team et al., 2024), Mixtral 8×7B (Jiang et al., 2024), LLaMA 3.3-70B (Dubey et al., 2024), and Qwen2.5 72B (Yang et al., 2024). These models span different architectures and parameter sizes, allowing for a comprehensive comparison on OCR text.
Traditionally, QA systems are evaluated using Exact Match (EM). However, these metrics can be insufficient for LLMs, as models often generate verbose responses, leading to low EM scores even when the correct answer is included in the response. To address this limitation, we evaluate MultiOCR-QA using BERTScore (Yuan et al., 2021) alongside EM. Additionally, we introduce another evaluation metric, Contains, to better assess MultiOCR-QA performance. Contains measures the extent to which the ground truth is present in the response generated by the model, regardless of the verbosity. We will apply these metrics to evaluate the QA results on RawOCR texts and then on CorrectedOCR texts used as context.
@@ 5.2. Experimental Results
English text: Table 3 presents the performance of LLMs on English text using both CorrectedOCR (CP) and RawOCR (RP) paragraphs. Across all models, the transition from CP to RP negatively impacts performance. The best-performing model on CP is LLaMA-3.3-70B with a BERTScore of 66.94, followed by Gemma-2 27B at 63.99. When switching to RP, LLaMA-3.3-70B still achieves the highest BERTScore 62.87, but with a 6.08% drop, highlighting its robustness. The lowest-performing model is Mixtral 8×22B, showing the most significant impact of OCR errors. LLaMA-3.3-70B model achieves the best performance in CP for the Contains metric of 63.90, indicating its strong retrieval ability in clean text. However, it experiences a 21.3% drop in RP, suggesting moderate sensitivity to OCR errors. The lowest performing model in RP is again Mixtral 8×22B for the Contains metric, which sees a 15.70% decrease, indicating its greater vulnerability to noisy text. The Exact Match (EM) metric shows the steepest decline across models, emphasizing that OCR errors severely impact the models’ ability to generate precise answers.
French text: In Table 4 we summarize how LLMs perform on QA over French text. Models consistently perform better on CorrectedOCR, confirming and quantifying the negative influence of OCR errors on the QA accuracy. Gemma-2 27B model achieves the highest BERTScore of 76.51 on CP, demonstrating its strong ability to capture semantic similarity. Despite the 1.46% drop, it still maintains the highest performance with a BERTScore of 75.39 on RP, indicating robustness to OCR noise. Mixtral 8×22B model shows the smallest drop of 1.05%, but its overall score remains lower than the ones for the other models. For Contains, Qwen-2.5 72B achieves the highest score on CP 57.55, highlighting superior retrieval performance on clean text. However, it experiences a 19.90% drop in RP, reinforcing its vulnerability to OCR errors. LLLaMA-3.1 8B also struggles, with a 19.63% decline, showing difficulty in retrieving spans from noisy text. In terms of EM, Gemma-2 27B outperforms all models with an EM of 17.46 on CP. Although it drops by 15.23%, it still maintains the best EM value 14.80 in RP. Mixtral 8×22B model is the most affected in EM, dropping by 27.99%, suggesting that OCR noise drastically reduces its accuracy in generating exact answers.
Table 3. Performance of LLMs on English Language: Comparison Using CorrectedOCR Paragraphs (CP) and RawOCR Paragraphs (RP) as Context. Red numbers indicate the percentage decrease in performance with RP. Bold values highlight the highest performance for each metric with CP, while underlined values denote the best performance for each metric with RP.
@@
Table 4. Performance of LLMs on French Language: Comparison Using CorrectedOCR Paragraphs (CP) and RawOCR Paragraphs (RP) as Context. Red numbers indicate the percentage decrease in performance with RP. Bold values highlight the highest performance for each metric with CP, while underlined values denote the best performance for each metric with RP.
@@
Table 5. Performance of LLMs on German Language: Comparison Using CorrectedOCR Paragraphs (CP) and RawOCR Paragraphs (RP) as Context. Red numbers indicate the percentage decrease in performance with RP. Bold values highlight the highest performance for each metric with CP, while underlined values denote the best performance for each metric with RP.
@@
German text: Table 5 focusing on German language shows the most significant performance decline among the three languages. Unlike English and French, German exhibits the largest performance drop due to its lower OCR quality, and the models struggle more with its linguistic structure. Among the models evaluated, Gemma-2 27B achieves the highest BERTScore of 67.07 on CP, confirming its strong ability to capture semantic similarity in clean text. It also maintains the best performance on RP (63.78); however, it still experiences a 4.91% decrease, highlighting the adverse effects of OCR errors. In contrast, LLaMA-3.1 8B and LLaMA-3.3 70B show the biggest BERTScore drop (8.74% and 8.43% respectively), suggesting that these models struggle more with OCR noise. For the Contains metric, LLaMA-3.3 70B achieves the highest Contains score on CP, making it the most effective for retrieving relevant information in clean text. However, all models suffer a severe drop in Contains when moving to RP, with reductions exceeding 65%. Qwen-2.5 7B has the worst drop 66.40%, indicating that it faces challenges to retrieve information from noisy text. In terms of EM, Gemma-2 27B achieves the highest EM score on CP 2.69, while Mixtral 8×22B has the lowest EM 0.078. EM scores drop drastically across all models, the largest decrease is 79.04% for Gemma-2, reinforcing that word-level distortions from OCR errors make exact answer matching nearly impossible. The Mixtral model drops 72.74% in EM, making it highly unreliable for exact answers in noisy OCR text.
Summary of findings: OCR errors consistently degrade the performance of the models in English, French and German texts, resulting in maximum drop of 9.23%, 1.75%, and 8.74% in BERTScore, respectively. The most severe impact was observed in German due to the lower quality of the OCR and the complex linguistic structure. While larger models like Gemma-2 27B and LLaMA-3.3 70B demonstrate greater resilience, all models suffer substantial declines in Contains and Exact Match (EM) metrics, highlighting their weakness in retrieving and generating precise answers from noisy text. Gemma-2 27B consistently outperforms others, maintaining the highest BERTScore and EM across all languages, but still experiences notable degradation in noisy conditions. Mixtral 8×22B emerges as the most vulnerable, exhibiting the lowest absolute performance and struggling particularly with exact answer generation.

### v3

@@ Abstract.
journalyear: 2025copyright: ccconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Koreabooktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreadoi: 10.1145/3746252.3761295isbn: 979-8-4007-2040-6/2025/11ccs: Information systems Question answeringccs: Information systems Content analysis and feature selection
@@ 1. Introduction
Historical corpora have already been used as an underlying text for numerous natural language processing (NLP) tasks, including Named Entity Recognition (NER), topic modeling, text classification, neural ranking, and understanding semantic changes in language over time (Ehrmann et al., 2020, 2023; Yang et al., 2011; Liebeskind and Liebeskind, 2020; Subramani et al., 2020). While these tasks offer valuable insights, it is expected that automatic question answering (QA) systems become an important interface to the vast historical collections. QA systems enable direct and intuitive access to information, offering a powerful means of analyzing, searching and understanding historical texts. By retrieving precise answers to user queries, QA can significantly enhance the accessibility, interpretability, and utility of historical corpora, making them more easily accessible and more useful for scholars, researchers, and the general public alike.
Figure 1. An example of CorrectedOCR and RawOCR text from the MultiOCR-QA dataset for the English language, highlighting different types of errors along with questions corresponding to this text. WER and CER denote Word Error Rate and Character Error Rate, respectively, indicating the level of errors in the text. The green highlights represent insertion errors, where IEO denotes Insertion Edit Operations - the number of insertions needed to transform RawOCR into CorrectedOCR. Similarly, red and blue highlights indicate deletion and substitution errors, with DEO and SEO representing Deletion Edit Operations and Substitution Edit Operations, respectively. The black boxes with numbers in the CorrectedOCR and RawOCR text correspond to the answers for each question in the paragraph.
@@
Although extensive research has focused on improving OCR accuracy and post-processing correction techniques, the specific impact of OCR noise on QA remains largely unexplored. Previous studies have examined challenges related to OCR in information retrieval (IR) (de Oliveira et al., 2023), historical text processing (Piotrowski, 2012; Hill and Hengchen, 2019), and entity recognition (Grover et al., 2008). However, a systematic investigation of how OCR errors affect QA model performance is still missing. Furthermore, our study contributes also to the body of work on robustness of LLMs against noisy inputs such as noisy prompts (Wang et al., 2023), however we do it in a quite novel yet realistic setting.
In this paper, we address this gap in understanding how LLMs perform in QA tasks when dealing with noisy OCR-generated text. In particular, we first introduce MultiOCR-QA  [^2] The dataset is available at https://github.com/DataScienceUIBK/MultiOCR-QA, a new multilingual QA dataset covering English, French, and German. This dataset includes both RawOCR text (OCR-generated text with errors) and CorrectedOCR text (ground truth), allowing for a direct comparison of QA performance under different text quality conditions. To generate contextually relevant question-answer pairs from historical text excerpts, we leverage instruction-fine-tuned LLMs. We then systematically evaluate the impact of different types of OCR errors—insertions, deletions, and substitutions—on QA model performance, offering new insights into the strengths and limitations of LLMs when processing noisy historical data. Figure 1 illustrates an example from the MultiOCR-QA dataset with various OCR errors impacting text accuracy. It presents QA pairs and the corresponding responses for questions when using CorrectedOCR and RawOCR text as input context to Gemma 2–27b, demonstrating the effects of OCR noise on QA performance. For instance, Question 1: ”What did the man say happened to his leg?” produces different answers depending on whether clean or noisy text was input. When using the CorrectedOCR context, the model produces the response: ”He said his sense of the leg was gone.” In contrast, the response based on the RawOCR context is: ”The fence of his leg is gone,” demonstrating how OCR-induced noise can significantly alter the meaning of the answer and degrade the QA accuracy.
@@ 3.1. Data Collection
  • French: For the French language, the ICDAR 2019 dataset provides a collection of 2,800 files obtained from three sources: the HIMANIS [^6] https://www.himanis.org Project, IMPACT - National Library of France, and the RECEIPT [^7] http://findit.univ-lr.fr/ dataset.
@@ 3.2. Question-Answer Generation
For fine-tuning the model in English, we used the SQuAD v1 dataset (Rajpurkar et al., 2016). We randomly selected 2,067 paragraphs and 10,570 questions from the development set and 3,000 paragraphs and 13,894 questions from the test set.
Similarly, for the French language, we fine-tuned the model on the FQuAD dataset (d’Hoffschmidt et al., 2020), utilizing both the validation and training sets, comprising 5,689 paragraphs and 23,919 questions. For the German language, we fine-tuned the model on the GermanQuAD dataset (Möller et al., 2021) using the training and validation sets, which consisted of 3,014 paragraphs and 13,722 questions. This fine-tuning step ensured that the model accurately generated factoid-style QA pairs, reducing instances of open-ended or conversational questions.
After instruction fine-tuning, we used the fine-tuned model for each language to generate questions for the preprocessed dataset prepared in the initial step. To generate high-quality QA pairs, we employed a two-step prompt-based approach: Answer Extraction and Question Generation for the Extracted Answers.
@@
Dataset Filtering: After generating the dataset for each language, we applied additional filtering steps to ensure quality and consistency. Specifically, we removed questions that did not end with a question mark, duplicate questions, and questions with excessively long answers. Since LLM-generated datasets may contain hallucinated long answers, we applied this additional filtering by removing excessively long answers. As a result, we removed 66 questions for English Language, 530 questions for French Language, and 7,210 questions for German Language. This final filtering step ensured that the MultiOCR-QA dataset consisted of concise, well-structured, and factually accurate question-answer pairs.
@@ 4. Dataset Analysis
After applying all the filtering steps, we obtained the final dataset, comprising 50,079 question-answer pairs. The dataset statistics, including average paragraph length, question length, and answer length, are presented in Table 1.
4.1. Quantifying and Filtering Noise
To assess the impact of OCR noise on QA quality, we quantified the noise level in the RawOCR text using two standard metrics: Character Error Rate (CER) and Word Error Rate (WER). CER measures the proportion of character-level errors in the RawOCR text compared to the ground truth text. It is computed as the number of insertions, deletions, and substitutions (including spaces) required to transform the RawOCR text into its correct form. WER quantifies word-level discrepancies, representing the proportion of words that require modifications (insertions, deletions, or substitutions) to match ground truth. Both CER and WER were computed using the Levenshtein distance (Miller et al., 2009), which determines the minimum number of edits needed to correct the OCR-generated text. A high CER but low WER suggests that errors are concentrated within a few words (e.g., spelling variations), whereas a high WER indicates distortions across multiple words, significantly affecting readability.
@@ 4.1. Quantifying and Filtering Noise
Avg. CorrectedOCR paragraph length (words) 271.73 297.53 217.33
@@
Avg. answer length (words) 2.93 3.12 5.63
@@
Outlier Detection: To ensure a reliable analysis, we applied the Interquartile Range (IQR) method to detect and remove outliers in CER values. Outliers were defined as CER values below Q1 - 1.5 * IQR or above Q3 + 1.5 * IQR, where Q1 and Q3 represent the 33rd and 66th percentiles of the CER distribution, respectively, and IQR is the difference between Q3 and Q1. This filtering resulted in the removal of 25 English, 351 French, and 2,423 German paragraphs.
@@ 5.1. Experimental Settings
We conducted experiments using multiple large language models (LLMs), including Qwen2.5 7B (Yang et al., 2024), LLaMa 3.1-8B (Dubey et al., 2024), Gemma-2-27B (Team et al., 2024), Mixtral 8x22B (Jiang et al., 2024), LLaMA 3.3-70B (Dubey et al., 2024), and Qwen2.5 72B (Yang et al., 2024). These models span different architectures and parameter sizes, allowing for a comprehensive comparison on OCR text.
Traditionally, QA systems are evaluated using Exact Match (EM). However, these metrics can be insufficient for LLMs, as models often generate verbose responses, leading to low EM scores even when the correct answer is included in the response. To address this limitation, we evaluate MultiOCR-QA using BERTScore (Yuan et al., 2021) alongside EM. Additionally, we introduce another evaluation metric, Contains, to better assess MultiOCR-QA performance. Contains measures the extent to which the ground truth is present in the response generated by the model, regardless of the verbosity.
We will apply these metrics to evaluate the QA results on RawOCR texts and then on CorrectedOCR texts used as context.
@@ 5.2. Experimental Results
English text: Table 3 presents the performance of LLMs on English text using both CorrectedOCR (CP) and RawOCR (RP) paragraphs. Across all models, the transition from CP to RP negatively impacts performance. The best-performing model on CP is LLaMA-3.3-70B with a BERTScore of 66.94, followed by Gemma-2 27B at 63.99. When switching to RP, LLaMA-3.3-70B still achieves the highest BERTScore 62.87, but with a 6.08% drop, highlighting its robustness. The lowest-performing model is Mixtral 8x22B, showing the most significant impact of OCR errors. LLaMA-3.3-70B model achieves the best performance in CP for the Contains metric of 63.90, indicating its strong retrieval ability in clean text. However, it experiences a 21.3% drop in RP, suggesting moderate sensitivity to OCR errors. The lowest performing model in RP is again Mixtral 8x22B for the Contains metric, which sees a 15.70% decrease, indicating its greater vulnerability to noisy text. The EM metric shows the steepest decline across models, emphasizing that OCR errors severely impact the models’ ability to generate precise answers.
French text: In Table 4 we summarize how LLMs perform on QA over French text. Models consistently perform better on CorrectedOCR, confirming and quantifying the negative influence of OCR errors on the QA accuracy. Gemma-2 27B model achieves the highest BERTScore of 76.51 on CP, demonstrating its strong ability to capture semantic similarity. Despite the 1.46% drop, it still maintains the highest performance with a BERTScore of 75.39 on RP, indicating robustness to OCR noise. Mixtral 8x22B model shows the smallest drop of 1.05%, but its overall score remains lower than the ones for the other models. For Contains, Qwen-2.5 72B achieves the highest score on CP 57.55, highlighting superior retrieval performance on clean text. However, it experiences a 19.90% drop in RP, reinforcing its vulnerability to OCR errors. LLaMA-3.1 8B also struggles, with a 19.63% decline, showing difficulty in retrieving spans from noisy text. In terms of EM, Gemma-2 27B outperforms all models with an EM of 17.46 on CP. Although it drops by 15.23%, it still maintains the best EM value 14.80 in RP. Mixtral 8x22B model is the most affected in EM, dropping by 27.99%, suggesting that OCR noise drastically reduces its accuracy in generating exact answers.
Table 3. Performance of LLMs on English Language: Comparison of CorrectedOCR (CP) and RawOCR Paragraphs (RP) as Context. Red numbers indicate the percentage decrease in performance with RP. Bold values highlight the highest performance for each metric with CP, while underlined values denote the best performance for each metric with RP.
@@
Table 4. Performance of LLMs on French Language: Comparison Using CorrectedOCR (CP) and RawOCR Paragraphs (RP) as Context. Red numbers indicate the percentage decrease in performance with RP. Bold values highlight the highest performance for each metric with CP, while underlined values denote the best performance for each metric with RP.
@@
Table 5. Performance of LLMs on German Language: Comparison Using CorrectedOCR (CP) and RawOCR Paragraphs (RP) as Context. Red numbers indicate the percentage decrease in performance with RP. Bold values highlight the highest performance for each metric with CP, while underlined values denote the best performance for each metric with RP.
@@
German text: Table 5 focusing on German language shows the most significant performance decline among the three languages. Unlike English and French, German exhibits the largest performance drop due to its lower OCR quality, and the models struggle more with its linguistic structure. Among the models evaluated, Gemma-2 27B achieves the highest BERTScore of 67.07 on CP, confirming its strong ability to capture semantic similarity in clean text. It also maintains the best performance on RP (63.78); however, it still experiences a 4.91% decrease, highlighting the adverse effects of OCR errors. In contrast, LLaMA-3.1 8B and LLaMA-3.3 70B show the biggest BERTScore drop (8.74% and 8.43% respectively), suggesting that these models struggle more with OCR noise. For the Contains metric, LLaMA-3.3 70B achieves the highest Contains score on CP, making it the most effective for retrieving relevant information in clean text. However, all models suffer a severe drop in Contains when moving to RP, with reductions exceeding 65%. Qwen-2.5 7B has the worst drop 66.40%, indicating that it faces challenges to retrieve information from noisy text. In terms of EM, Gemma-2 27B achieves the highest EM score on CP 2.69, while Mixtral 8x22B has the lowest EM 0.078. EM scores drop drastically across all models, the largest decrease is 79.04% for Gemma-2, reinforcing that word-level distortions from OCR errors make exact answer matching nearly impossible. The Mixtral model drops 72.74% in EM, making it highly unreliable for exact answers in noisy OCR text.
Summary of findings: OCR errors consistently degrade the performance of the models in English, French and German texts, resulting in maximum drop of 9.23%, 1.75%, and 8.74% in BERTScore, respectively. The most severe impact was observed in German due to the lower quality of the OCR and the complex linguistic structure. While larger models like Gemma-2 27B and LLaMA-3.3 70B demonstrate greater resilience, all models suffer substantial declines in Contains and Exact Match (EM) metrics, highlighting their weakness in retrieving and generating precise answers from noisy text. Gemma-2 27B consistently outperforms others, maintaining the highest BERTScore and EM across all languages, but still experiences notable degradation in noisy conditions. Mixtral 8x22B emerges as the most vulnerable, exhibiting the lowest performance and struggling particularly with exact answer generation.